| TIP: Click on subject to list as thread! | ANSI |
| echo: | |
|---|---|
| to: | |
| from: | |
| date: | |
| subject: | 4\22 Pt 1 ISS On Orbit Status 22-04-2003 |
This Echo is READ ONLY ! NO Un-Authorized Messages Please!
~~~~~~~~~~~~~~~~~~~~~~~~ ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~
23 April 2003
ISS On-Orbit Status 22 Apr 2003
Part 1 of 2
All ISS systems continue to function nominally, except as noted
previously or below. Day 150 for Expedition 6 (since STS-113 launch,
11/23/02).
FE-1 Nikolai Budarin performed the third experiment run of the
Russian Plasma Crystal-3 (PK-3, Plazmennyi Kristall 3) payload, after
activating the turbopump for evacuating the work chamber and setting
up for video recording. The pump will be turned off in the
evening. [During the day, Budarin monitored the experiment, which
involved, first, warming up the ZB experiment unit for at least 3
hrs, followed by three parts: (1) studying particle agglomeration
(no plasma, 25 min); (2) studying isolated negative pressure wave
distribution in structures (20 min); and (3) studying the residual
charge of minute particles after charge deactivation (relaxation)
under varying pressures (40 min).]
The photos of the disassembled ARCTIC-1 refrigerator/freezer
equipment taken by FE-2/SO Don Pettit yesterday are being analyzed.
Based on yesterday's discussion, instructions were uplinked for
Pettit to attempt to bypass/work around failed TEDs (thermoelectric
devices) with jumpers which must be soldered in place. Today's
schedule had 2h 15m reserved for this IFM. [Each of the TEDs (i.e.,
the heat pumps providing the refrigeration function) consists of
many (~100) semiconductor cells (crystalline Bismuth Telluride,
Bi2Te3) in series, arranged in two stages in form of a double-decker
sandwich between aluminum oxide ceramic plates. If the two stages
are no longer correctly connected, possibly due to corrosion, new
solder junctions are required.]
CDR Kenneth Bowersox completed the periodic inspection and cleaning
of bacteria filters and smoke detectors (SDs), today in the Node (Lab
and Joint Airlock were done Sunday, 4/20), using pressurized N2
(nitrogen) gas for SD cleaning. The activities were video-recorded
and the video equipment stowed after the inspection/cleaning. [Last
time done: 11/18/02, by Peggy Whitson.]
Budarin worked on the Service Module (SM) plumbing, today removing
and replacing the gas-liquid mixture filter (FGS) as well as the
filter-reactor unit (F-R) in the powered-down SRV-K condensate water
processor. Both filters were then prepared for disposal on
Progress. [FGS was replaced last on 11/5/02 by Sergei Treschev.]
After the R&R, Budarin installed a new condensate separator and
sample container in the SKV-K2M upstream of the FGS. The SKV-K was
then turned on again, and later in the day (4:40pm EDT), Nikolai is
to return to the SKV to check behind a wall panel to make sure that
the SKV1 fan (VR) is not spreading moisture.
Don Pettit performed the daily routine maintenance of the SOZh life
support system (including the ASU toilet facility) and prepared the
daily IMS (inventory management system) "delta" file for updating the
IMS database.
Nikolai completed his regular daily 5-min. inspection of the BIO-5
Rasteniya-2/Lada-2 ("Plants-2") plant growth experiment.
The crew continued their preparations for the upcoming handover
activities to the Expedition 7 crew on 4/28-5/3 and their own return
on Soyuz TMA-1 on 5/4.
All crewmembers worked out on the RED resistive trainer, CEVIS cycle
ergometer and TVIS treadmill, completing their daily physical
exercise program. On the treadmill in motorized mode, Sox is limited
to 5.5 mph maximum speed, Nikolai to 7 mph and Don to 6 mph. Budarin
and Pettit also performed the weekly TVIS maintenance.
Shortly before dinner time (3:45pm), Nikolai Mikhailovich will break
out and set up the equipment required for the Russian MedOps
Reflotron biochemical blood test (MO-11), consisting of the Reflotron
IV analyzer, with accessories, power supply and Reflotron IV kit.
[Budarin will undergo the tests tomorrow (4/23), preceded by imbibing
250 ml of warm water or plain (unsugared) tea 20 minutes before
taking the blood samples.]
At 12:10am EDT, the three crewmembers engaged in an excellent
PAO/educational TV Q&A session with students and teachers assembled
at the Gulf Coast Exploreum Science Center in Mobile, Alabama.
["Cosmonaut Budarin, as the only Russian on the station, is your work
different from the work the Americans do?"; "On a given day when I
look outside my window, I might see a neighbor walking her dog or
kids playing basketball. Can you give a description of what you see
when you look out your window? "]
The successful checkout of the cardiac defibrillator by the CDR last
week earned words of praise from the Flight Surgeon. As a result,
the defib can be left in its current configuration, plugged in with
the PDIM (power data interface module) off and the batteries stowed
for their protection.
Don is scheduled today to discuss possible additional InSPACE
(Investigating the Structure of Paramagnetic Aggregates from
Colloidal Emulsions) payload operations in the MSG (Microgravity
Science Glovebox) in a telecon with the experiment's Principal
Investigator (PI) from MIT.
Because of the Russian MBI-4 "Farma" drug assessment session
scheduled for 4/24, Budarin was advised to suspend ingestion of all
pharmacological preparations from today through Thursday.
The R3 software upgrade of the two EXT MDM (external systems
computers) was successfully accomplished, without any hitches in the
transition. Checksums are nominal. [The step-up is required to
support P3 and P4 truss installation on Mission 12A operations. The
new functionality remains inhibited until truss activation procedures
on 12A.]
Launch preparations for 6S/Soyuz at the Baikonur Cosmodrom remain on
schedule. Rollout of the Soyuz launcher plus spacecraft will take
place next Thursday (4/24). Launch of 6S is currently still set for
11:54pm EDT on 4/25 (Friday night).
Today's CEO (crew earth observations) targets, currently restricted
by limited the use of the science window in the Lab including
near-vertical targets from the city target list due to the current
XPOP attitude, were Rome, Italy (nadir pass; ESC [electronic still
camera]), Havana, Cuba (nadir pass; ESC), Monterrey, Mexico (nadir
pass; ESC), and Northern Europe (Dynamic event. Weather predicted
clear from Britain to Germany. Suggested panoramic views north and
south of track). CEO images can be viewed at the websites
http://eol.jsc.nasa.gov and http://earthobservatory.nasa.gov
- Continued -
@Message posted automagically by IMTHINGS POST 1.30
---
* Origin: SpaceBase(tm) Pt 1 -14.4- Van BC Canada 604-473-9358 (1:153/719.1)SEEN-BY: 633/267 270 @PATH: 153/719 715 7715 140/1 106/2000 633/267 |
|
| SOURCE: echomail via fidonet.ozzmosis.com | |
Email questions or comments to sysop@ipingthereforeiam.com
All parts of this website painstakingly hand-crafted in the U.S.A.!
IPTIA BBS/MUD/Terminal/Game Server List, © 2025 IPTIA Consulting™.